# The Architecture and Evaluation of Bayesian Neural Networks

### v1

@@ Abstract
As modern neural networks get more complex, specifying a model with high predictive performance and sound uncertainty quantification becomes a more challenging task. Despite some promising theoretical results on the true posterior predictive distribution of Bayesian neural networks, the properties of even the most commonly used posterior approximations are often questioned. Computational burdens and intractable posteriors expose miscalibrated Bayesian neural networks to poor accuracy and unreliable uncertainty estimates. Approximate Bayesian inference aims to replace unknown and intractable posterior distributions with some simpler but feasible distributions. The dimensions of modern deep models coupled with the lack of identifiability make Markov chain Monte Carlo tremendously expensive and unable to fully explore the multimodal posterior. On the other hand, variational inference benefits from improved computational complexity but lacks the asymptotical guarantees of sampling-based inference and tends to concentrate around a single mode. The performance of both approaches heavily depends on architectural choices; this paper aims to shed some light on this, by considering the computational costs, accuracy and uncertainty quantification in different scenarios including large width and out-of-sample data. To improve posterior exploration, different model averaging and ensembling techniques are studied, along with their benefits on predictive performance. In our experiments, variational inference overall provided better uncertainty quantification than Markov chain Monte Carlo; further, stacking and ensembles of variational approximations provided comparable to Markov chain Monte Carlo accuracy at a much-reduced cost.
Keywords: Bayesian Deep Learning, Approximate Bayesian Inference.
@@ 1 Introduction
Despite the tremendous success of deep learning in areas such as natural language processing [Touvron et al., 2023] and computer vision [Krizhevsky et al., 2017, Dosovitskiy et al., 2020], often there is no clear understanding of why a particular model performs well [Zhang et al., 2021, Szegedy et al., 2014]. Even though the universal approximation theorem guarantees that a wide enough feed-forward neural network with a single hidden layer can express any smooth function [Hornik et al., 1989], in practice, constructing a model which is not only expressive but generalizes well is challenging. In contrast, the so-called no free lunch theorem [Wolpert, 1996] dictates that there is no panacea to solve every problem, and one should be careful when designing a model appropriate to the task. Many modern machine learning models are over-parametrized and prone to overfitting, especially given the limited size of the dataset. Complex problems demand exploring bigger model spaces, and there is a danger of choosing an excessively over-parametrized model which is going to overfit and has a high variance. Additionally, conventional deep models do not offer human-understandable explanations and lack interpretability [Lipton, 2018]. By default, classical neural networks do not address the uncertainty associated with their parameters and whilst there exist proposals enabling neural networks (NNs) to provide some uncertainty estimates, they are often miscalibrated [Guo et al., 2017]. As a result, these models are typically overconfident, provide a low level of uncertainty even when data variations occur [Ovadia et al., 2019], and are easily fooled and are susceptible to adversarial attacks [Szegedy et al., 2014, Nguyen et al., 2015]. At the same time, reliable uncertainty quantification (UQ) is crucial for any decision-making process, and it is not enough to obtain a point estimate of the prediction. The key distinguishing property of the Bayesian framework is that it incorporates domain expertise and deals with uncertainty quantification in a principled way: by marginalizing over the posterior distribution of parameters. As a result, Bayesian models are more resistant to distribution shifts and can improve the accuracy and calibration of classical deep models [Wilson and Izmailov, 2020]. Nevertheless, the reliability of uncertainty estimates and the gap within-the-sample and out-of-sample performance still require improvement [Foong et al., 2020]. The distributions arising in Bayesian neural networks (BNNs) are typically infeasible and highly multimodal, and the core challenge lies in estimating the posterior [Papamarkou et al., 2022]. One should not only find a model that matches the task but, as importantly, achieve the alignment between the model and the applied inference algorithm [Gelman et al., 2020]; and the most theoretically grounded sampling methods and approximation techniques are limited by the computing budget, size of the dataset, and sheer number of parameters. We list several characteristics of classical and Bayesian neural networks in the Table 1.
@@
[…]
@@
Outline. In this work, we consider some of the challenges and nuances of Bayesian neural networks and evaluate the performance with different architectures and for different posterior inference algorithm choices. Specifically, we study the sensitivity of BNNs to the choice of width in Section 2.3, depth in Section 2.4, and investigate the performance of BNN under the distribution shift in Section 2.5; Across all the experiments in Section 2, we observe that for different inference algorithms, one model can provide strikingly diverse performances. The challenge of comparative model assessment is addressed in Section 3.1, where we introduce the estimated pointwise loglikelihood as a measure of model utility. While, given some set of models, the Bayesian approach has the potential to deal with the model choice by comparing posterior model probabilities, such comparison tends to favour one candidate disproportionally strongly [Oelrich et al., 2020]. Thus, the classical Bayesian model averaging (BMA) based on model probabilities [Hoeting et al., 1999] is only optimal if the true model is among the comparison set. In response to the limitations of BMA, in Sections 3.2 and 3.3 we consider ensembling, stacking and pseudo-BMA[Yao et al., 2018].
@@ 2.1 Architecture Components
Whilst the dimensions of the input and the output are determined by the dimensionality of the data set, the dimension of the weight space plays an essential part in specifying neural networks and can be tuned to improve prediction performance. In the case of feed-forward neural networks, this amounts to finding optimal depth and width. While the universal approximation theorem advocates for single-layer neural networks, deep neural networks gained popularity due to their expressiveness and tremendous success in real-world applications allowed by the increase in available computing power [Chatziafratis et al., 2020]. At the same time, the more parameters one has, the more nuanced the choice of the model becomes. No matter what the prediction task is, overly complex models suffer from the curse of dimensionality which causes not only poor performance but also computational problems. On a slightly different line, we recall the seminal result first obtained for neural networks with one hidden layer [Neal, 1995] and then extended to arbitrary depth [Matthews et al., 2018] which states that as the width of BNN tends to infinity, the distribution of the network’s output induced by the prior converges to the Gaussian process (GP) with a neural network kernel, also known as the neural network GP (NNGP); there is a similar correspondence relating GPs and distributions induced by the posterior [Hron et al., 2022]. When defining Bayesian neural networks, choosing a prior and understanding how properties and prior beliefs on the weight space translate to the functions is a major challenge. Even though we acknowledge the importance of prior specification, we do not empirically study different choices here. Note that generally, we require priors which are: (1) interpretable, e.g. we want to be able to specify the hyperparameters of the prior based on the task at hand; (2) have large support, i.e. prior should not concentrate around a small subset of the parameter space; (3) lead to feasible inference and favour reasonable approximations of the posterior and predictive distributions.
Finally, to specify any neural network one needs to choose the activation function which (apart from being nonlinear) is required to be differentiable. In our experiments, we consider a popular nowadays rectified linear unit function (ReLU) defined as $\max(0,x)$, which switches the negative inputs off and leaves the positive ones unchanged, as well as the sigmoid activation function defined as $\sigma(x)=\exp(x)/(\exp(x)+1)$.
2.2 Settings of the Experiment
In the experiment, we consider the following Bayesian neural network, illustrated by the Figure 1
Figure 1: Example of the directed acyclic graph (DAG) of the neural network used in the experiments when $L=2$.
@@ 2.2 Settings of the Experiment
where we consider two different choices of nonlinearity $g$, namely, the ReLU and the sigmoid and the following priors on the weights and biases:
$\bm{W}_{l}\sim\text{N}(0,\frac{1}{\sqrt{LD_{0}}})\quad\bm{W}_{l}\sim\text{N}(0,\frac{2}{\sqrt{D_{l-1}}}),\text{ for }l=2,\ldots,L+1,$
$\bm{b}_{l}\sim\text{N}(0,\frac{1}{2\sqrt{L}})\text{ for }l=1,\ldots,L+1,$
@@
2.3 Limiting the Width of the Network
We consider a simple synthetic dataset with one-dimensional input and output $\sin(10\bm{x})\bm{x}^{2}$, where the noisy observations on which the neural network is trained are obtained by adding some Gaussian noise:
@@ 2.3 Limiting the Width of the Network
The input is scaled, unlike the output. The training data $\bm{\mathcal{D}}$ consists of $N=500$ observations and the new data for testing $\tilde{\bm{\mathcal{D}}}$ consists of $\tilde{N}=100$ observations. We first study the performances of mfVIR, mfVIS, HMCR and HMCS with 1 hidden layer as the width increases and illustrate the metrics for $D_{1}=20,200,1000$ and $2000$ hidden units by the Figure 2(a). The predictions of all four models when $D_{1}=2000$ are provided on the Figure 2(b). The performance of the mfVIS dips with the increase in the dimension of the hidden layer, moreover, for $D_{1}=1000$ and $D_{1}=2000$ its posterior predictive distribution fails to capture the data, and, in fact, degenerates to the prior (Figure 2(b)). An explanation of why such behaviour occurs was obtained via the correspondence of Gaussian processes and BNNs. While as the width increases the true posterior of a BNN converges to NNGP posterior [Hron et al., 2022], any optimal mean-field Gaussian variational posterior of a BNN with odd (up to a constant offset) Lipschitz activation function converges to the prior predictive distribution of the NNGP [Coker et al., 2022]. In other words, the mean-field variational approximations of wide BNNs with sigmoid activation function ignore the data. If one abandons the mean-field assumption and proposes a full-rank variational family, then using variational inference (VI) for wider networks would take at least a hundred times more time than using HMC, which undermines the benefits of using VI. Such degenerate behaviour is not observed with HMC,(Figure 2(b)), but this comes at a subsequent significant increase in training time. For wider networks, the HMCR model exhibits a better performance than the HMCS both in terms of accuracy and uncertainty quantification. In terms of predictive accuracy, HMC is preferred over mfVI in all of the combinations of the activation function and width. However, in terms of uncertainty quantification, the HMC is inferior to mfVI. In our experiment, HMC underestimates the uncertainty of the signal much more than VI (Figures 2(b) and 2(a)). Note that whilst variational inference is often cursed to underestimate the uncertainty[Trippe and Turner, 2018], that is not always the case [Blundell et al., 2015, Gal and Ghahramani, 2016]. Markov chain Monte Carlo (MCMC) methods are known to struggle to effectively explore multimodal posteriors [Papamarkou et al., 2022, Izmailov et al., 2021], and lack of uncertainty could be a result of poor mixing of the chain.
@@
(a) Performance of models as the number of hidden units increases.
(b) The predictions and uncertainty estimates obtained by each model when $D_{1}=2000$
Figure 2: Predictive performance of wider BNNs.
2.4 Deeper Networks
Consider the data of Section 2.3 and neural networks defined by Equation 1 with the number of layers $L$ varying from 1 to 6 and a fixed number of hidden units in each layer $D_{h}=20$. Figure 3(a) provides the recorded metrics, and Figure 3(b) illustrates the predictions of all the four models with $L=6$. First, observe that overall both RMSE and empirical coverage of mfVIR approximations improve with the increase of depth. The mfVIS follows a similar pattern, except for the $L=5$ when the prediction quality of the network drops drastically. Indeed, the approximate posteriors of deep neural networks obtained with the mean-field variational inference were shown to be as flexible as the much richer approximate posteriors of shallower BNNs [Farquhar et al., 2020]. We do not obtain the same improvement in the prediction quality of models trained with HMC: the performance of HMCR falls whilst the HMCS does not improve as the depth increases. This undesirable behaviour could be a result of the multimodality of distributions in overparametrized models combined with the challenges of MCMC in exploring the high-dimensional space[Izmailov et al., 2021, Papamarkou et al., 2022]. Compared to the findings of Section 2.3, we note that the deeper NNs are less sensitive to the choice of the activation function. It is needless to say that the HMC algorithm scales rather poorly and as the number of layers changes from $L=1$ to $L=6$ the time needed to train HMCR and HMCS gets more than 15 and 30 times greater, respectively. We note that for models with more than one hidden layer, the network with sigmoid activations takes roughly twice as much time as the network with ReLU. The striking discrepancy in training times could arise due to the difference in the leapfrog integrator step sizes [Betancourt et al., 2015].
(a) Performance of models as the number of hidden layers increases.
(b) The predictions and uncertainty estimates obtained by each model when $L=6$
Figure 3: Prediction performance of deeper networks.
General summary. In terms of the training time, HMC becomes less and less feasible with the increase in depth. With the need to explore high-dimensional parameter spaces, multimodality of the posteriors should be kept in mind as an arising challenge for both mfVI and HMC. In terms of the balance between accuracy and UQ, the mean-field variational inference with ReLU activation function is able to outperform the MCMC with the increase in depth.

### v2

# Understanding the Trade-offs in Accuracy and Uncertainty Quantification: Architecture and Inference Choices in Bayesian Neural Networks

@@ Abstract
As modern neural networks get more complex, specifying a model with high predictive performance and sound uncertainty quantification becomes a more challenging task. Despite some promising theoretical results on the true posterior predictive distribution of Bayesian neural networks, the properties of even the most commonly used posterior approximations are often questioned. Computational burdens and intractable posteriors expose miscalibrated Bayesian neural networks to poor accuracy and unreliable uncertainty estimates. Approximate Bayesian inference aims to replace unknown and intractable posterior distributions with some simpler but feasible distributions. The dimensions of modern deep models, coupled with the lack of identifiability, make Markov chain Monte Carlo (MCMC) tremendously expensive and unable to fully explore the multimodal posterior. On the other hand, variational inference benefits from improved computational complexity but lacks the asymptotical guarantees of sampling-based inference and tends to concentrate around a single mode. The performance of both approaches heavily depends on architectural choices; this paper aims to shed some light on this, by considering the computational costs, accuracy and uncertainty quantification in different scenarios including large width and out-of-sample data. To improve posterior exploration, different model averaging and ensembling techniques are studied, along with their benefits on predictive performance. In our experiments, variational inference overall provided better uncertainty quantification than MCMC; further, stacking and ensembles of variational approximations provided comparable accuracy to MCMC at a much-reduced cost.
Keywords: Approximate Bayesian Inference, Bayesian Deep Learning, Ensembles, Out-of-Distribution, Uncertainty Quantification.
@@ 1 Introduction
Despite the tremendous success of deep learning in areas such as natural language processing [Touvron et al., 2023] and computer vision [Krizhevsky et al., 2017, Dosovitskiy et al., 2020], often there is no clear understanding of why a particular model performs well [Zhang et al., 2021, Szegedy et al., 2014]. Even though the universal approximation theorem guarantees that a wide enough feed-forward neural network with a single hidden layer can express any smooth function [Hornik et al., 1989], in practice, constructing a model which is not only expressive but generalizes well is challenging. In contrast, the so-called no free lunch theorem [Wolpert, 1996] dictates that there is no panacea to solve every problem, and one should be careful when designing a model appropriate to the task. Many modern machine learning models are over-parametrized and prone to overfitting, especially given the limited size of the dataset. Complex problems demand exploring bigger model spaces, and there is a danger of choosing an excessively over-parametrized model, which is going to overfit and have a high variance. Additionally, conventional deep models do not offer human-understandable explanations and lack interpretability [Lipton, 2018]. By default, classical neural networks do not address the uncertainty associated with their parameters and whilst there exist proposals enabling neural networks (NNs) to provide some uncertainty estimates, they are often miscalibrated [Guo et al., 2017]. As a result, these models are typically overconfident, provide a low level of uncertainty even when data variations occur [Ovadia et al., 2019], and are easily fooled and are susceptible to adversarial attacks [Szegedy et al., 2014, Nguyen et al., 2015]. At the same time, reliable uncertainty quantification (UQ) is crucial for any decision-making process, and it is not enough to obtain a point estimate of the prediction.
The key distinguishing property of the Bayesian framework is that it incorporates domain expertise and deals with uncertainty quantification in a principled way: by marginalizing with respect to the posterior distribution of parameters. As a result, Bayesian models are more resistant to distribution shifts and can improve the accuracy and calibration of classical deep models [Wilson and Izmailov, 2020]. Nevertheless, the reliability of uncertainty estimates and the gap between within-the-sample and out-of-sample performance still require improvement [Foong et al., 2020]. The posterior distributions arising in Bayesian neural networks (BNNs) are analytically unavailable and highly multimodal, and the core challenge lies in estimating the posterior [Papamarkou et al., 2022]. One should not only find a model that matches the task but, as importantly, achieve the alignment between the model and the applied inference algorithm [Gelman et al., 2020]; and the most theoretically grounded sampling methods and approximation techniques are limited by the computing budget, size of the dataset, and sheer number of parameters. We list several characteristics of classical and Bayesian neural networks in the Table 1.
@@
Interpretability poor improved $\checkmark$
Robustness to OOD poor improved $\checkmark$
Adversarial attacks sensitive less sensitive $\checkmark$
Overconfidence typical less typical $\checkmark$
@@
Outline. In this work, we consider some of the challenges and nuances of Bayesian neural networks and evaluate the performance with different architectures and for different posterior inference algorithm choices. Specifically, we study the sensitivity of BNNs to the choice of width in Section 2.3, depth in Section 2.4, and investigate the performance of BNNs under the distribution shift in Section 2.5. Across all the experiments in Section 2, we observe that for different inference algorithms, one model can provide strikingly diverse performances. The challenge of comparative model assessment is addressed in Section 3.1, where we introduce the estimated pointwise loglikelihood as a measure of model utility. While given some set of models, the Bayesian approach has the potential to deal with the model choice by comparing posterior model probabilities, such comparison tends to favour one candidate disproportionally strongly [Oelrich et al., 2020]. Thus, the classical Bayesian model averaging (BMA) based on model probabilities [Hoeting et al., 1999] is only optimal if the true model is among the comparison set. In response to the limitations of BMA, in Sections 3.2, 3.3 and 3.4 we consider ensembling, stacking and pseudo-BMA[Yao et al., 2018].
@@ 2.1 Architecture Components
Whilst the dimensions of the input and the output are determined by the dimensionality of the data set, the dimension of the weight space plays an essential part in specifying neural networks and can be tuned to improve prediction performance. In the case of feed-forward neural networks, this amounts to finding optimal depth and width. While the universal approximation theorem advocates for single-layer neural networks [Hornik et al., 1989], variants of the universal approximation theorem exist for deeper networks [Lu et al., 2017, Hanin, 2019]. Further, deep neural networks gained popularity due to their expressiveness and tremendous success in real-world applications, allowed by the increase in available computing power [Chatziafratis et al., 2020]. At the same time, the more parameters one has, the more nuanced the choice of the model becomes. No matter what the prediction task is, overly complex models suffer from the curse of dimensionality which causes not only poor performance but also computational problems.
On a slightly different line, we recall the seminal result first obtained for neural networks with one hidden layer [Neal, 1995] and then extended to arbitrary depth [Matthews et al., 2018] which states that under general conditions, as the width of a BNN tends to infinity, the distribution of the network’s output induced by the prior converges to a Gaussian process (GP) with a neural network kernel, also known as the neural network GP (NNGP); there is a similar correspondence relating GPs and distributions induced by the posterior [Hron et al., 2022]. When defining BNNs, choosing a prior and understanding how properties and prior beliefs on the weight space translate to the functions is a major challenge. Note that generally, we require priors which are: (1) interpretable, e.g. we want to be able to specify the hyperparameters of the prior based on the task at hand; (2) have large support, i.e. prior should not concentrate around a small subset of the parameter space; (3) lead to feasible inference and favour reasonable approximations of the posterior and predictive distributions.
Finally, to specify any neural network, one needs to choose the activation function, which (apart from being nonlinear) is required to be differentiable. In our experiments, we consider the widely-used rectified linear unit function (ReLU) defined as $\max(0,x)$, which switches the negative inputs off and leaves the positive ones unchanged, as well as the sigmoid activation function defined as $\sigma(x)=\exp(x)/(\exp(x)+1)$.
2.2 Setup of the Experiments
In the experiments, we consider the following BNN, illustrated by the Figure 1(a)
(a) Example of the DAG of the neural network when $L=2$.
(b) Examples of priors of $\bm{W}_{l}$ for $l=2,\ldots,L+1$ and $D_{l}=20$.
Figure 1: Examples of the directed acyclic graph (DAG) of the neural network and of the priors used in the experiments.
@@ 2.2 Setup of the Experiments
$\bm{W}_{1}\sim F\left(0,\frac{1}{4L}\right),\quad\bm{W}_{l}\sim F\left(0,\frac{4}{D_{l-1}}\right)\quad\text{for}\quad l=2,\ldots,L+1,$
$\bm{b}_{l}\sim\text{N}\left(0,\frac{1}{4L}\right)\quad\text{for}\quad l=1,\ldots,L+1,$
where the notation $F(\mu,\sigma^{2})$ represents a distribution with mean $\mu$ and scale $\sigma$, and specifically, here is chosen as either Gaussian or Student-t with 5 degrees of freedom. To avoid divergence in wider networks and mitigate the damage caused by the nonlinear deformation [He et al., 2015], the weights’ variance is scaled by the inverse of the preceding layer’s width.
The BNN defined by Equation 1 and trained with automatic differentiation variational inference (ADVI) [Kucukelbir et al., 2017], which assumes a mean-field (diagonal) Gaussian variational family, is referred to as mfVIR or mfVIS, depending on the choice of the activation: ReLU or sigmoid, respectively. The model trained with the Hamiltonian Monte Carlo (HMC), using the No U-Turn Sampler (NUTS) [Hoffman and Gelman, 2014] is denoted as HMCR or HMCS. For simplicity, we often refer to one-layer neural networks of particular width $D$ as to mfVIRD, mfVISD, HMCRD or HMCSD (e.g. one-layer BNN with 20 hidden units and ReLU activation trained with mean-field VI is called mfVIR20). All experiments are implemented with Numpyro [Phan et al., 2019], ArviZ [Kumar et al., 2019], JAX [Bradbury et al., 2018] and Flax [Heek et al., 2024]. We record the run time of the approximate inference (TT), the root mean squared error RMSE and empirical coverage for the function and observations (EC). Note that we compute empirical coverage as a fraction of observations contained within the $95\%$ confidence interval (CI), this means that in the ideal settings the computed EC should equal to $0.95$. If $\text{EC}>0.95$ then the confidence intervals are too wide; a worse scenario occurs when $\text{EC}<0.95$ as it means that the CIs are too narrow and the model is overconfident in predictions. Details on the computed metrics and the corresponding formulas are discussed in Appendix A, where we provide further information on the initialization and parameters for the inference algorithms [^1] The code is available on GitHub..
The absence of the test log-likelihood among the recorded metrics is motivated by the observation that the higher test log-likelihood does not necessarily correspond to a more accurate posterior approximation nor to lower predictive error (such as RMSE)[Deshpande et al., 2024].
2.3 Increasing the Width of the Network
We consider a simple synthetic dataset with one-dimensional input and output:
@@ 2.3 Increasing the Width of the Network
The training data $\bm{\mathcal{D}}$ consists of $N=500$ observations and the new data for testing $\tilde{\bm{\mathcal{D}}}$ consists of $\tilde{N}=100$ observations. We first study the performances of mfVIR, mfVIS, HMCR and HMCS with 1 hidden layer and either Gaussian or Student-t priors as the width increases, and illustrate the metrics for $D_{1}=20,200,1000$ and $2000$ hidden units by the Figure 2(a). The predictions of the four combinations of activation and inference algorithm with Gaussian priors when $D_{1}=2000$ are provided on the Figure 2(b); similar results were obtained when weights have Student-t distribution, the figures are presented in Appendix C. For either choice of priors, performance of the mfVIS dips with the increase in the dimension of the hidden layer; moreover, for $D_{1}=1000$ and $D_{1}=2000$ its posterior predictive distribution fails to capture the data, and, in fact, degenerates to the prior (Figure 2(b)). An explanation of why such behaviour occurs was obtained via the correspondence of Gaussian processes and BNNs. While as the width increases the true posterior of a BNN converges to a NNGP posterior [Hron et al., 2022], any optimal mean-field Gaussian variational posterior of a BNN with odd (up to a constant offset) Lipschitz activation function converges to the prior predictive distribution of the NNGP [Coker et al., 2022]. In other words, the mean-field variational approximations of wide BNNs with sigmoid activations ignore the data. If one abandons the mean-field assumption and proposes a full-rank variational family, then using variational inference (VI) for wider networks would take at least a hundred times more time than using HMC, which undermines the benefits of using VI. Such degenerate behaviour is not observed with HMC(Figure 2(b), but this comes at a significant increase in training time. For wider networks, the HMCR model exhibits a better performance than the HMCS both in terms of accuracy and uncertainty quantification. In terms of predictive accuracy, HMC is preferred over mfVI in all of the combinations of the activation function and width. However, in terms of uncertainty quantification, the HMC is inferior to mfVI (with one exception of a BNN with Student-t priors, sigmoid activation and 2000 hidden units). In our experiment, HMC underestimates the uncertainty of the signal much more than VI (Figures 2(b) and 2(a)). Note that whilst variational inference is often cursed to underestimate the uncertainty[Trippe and Turner, 2018], that is not always the case [Blundell et al., 2015, Gal and Ghahramani, 2016]. Markov chain Monte Carlo (MCMC) methods are known to struggle to effectively explore multimodal posteriors [Papamarkou et al., 2022, Izmailov et al., 2021], and a lack of uncertainty could be a result of poor mixing of the chain.
@@
(a) Metrics of methods as the number of hidden units increases.
(b) Predictions and uncertainty estimates for each method with $D_{1}=2000$ and Gaussian priors.
Figure 2: Predictive performance of BNNs as the width increases.
2.4 Increasing the Depth of the Networks
Consider the data of Section 2.3 and neural networks defined by Equation 1 with the number of layers $L$ varying from 1 to 6 and a fixed number of hidden units in each layer $D_{h}=20$. Figure 3(a) provides the recorded metrics, and Figure 3(b) illustrates the predictions of the four combinations of activation and inference algorithm with $L=6$ and Gaussian priors (analogous figures for Student-t priors are provided in Appendix C). First, observe that overall both RMSE and empirical coverage of mfVIR approximations improve with the increase of depth, with one exception of $L=5$ and Student-t priors, when the prediction quality of the network drops drastically. The mfVIS follows a similar pattern, except for the case of $L=5$ and Gaussian priors. Indeed, the approximate posteriors of deep neural networks obtained with the mean-field variational inference were shown to be as flexible as the much richer approximate posteriors of shallower BNNs [Farquhar et al., 2020]. We do not obtain the same improvement in the prediction quality of models trained with HMC: for either choice of priors, the performance of HMCR falls, whilst the HMCS does not improve as the depth increases. This undesirable behaviour could be a result of the multimodality of distributions in overparametrized models combined with the challenges of MCMC in exploring the high-dimensional space [Izmailov et al., 2021, Papamarkou et al., 2022]. Compared to the findings of Section 2.3, we note that the deeper NNs are less sensitive to the choice of the activation function. It is needless to say that the HMC algorithm scales rather poorly, and as the number of layers changes from $L=1$ to $L=6$, the time needed to train HMCR and HMCS gets more than 15 and 30 times greater, respectively. We note that for models with more than one hidden layer, training of the network with sigmoid activations takes roughly twice as much time as the network with ReLU. The striking discrepancy in training times could arise due to the difference in the leapfrog integrator step sizes [Betancourt et al., 2015].
(a) Metrics of methods as the number of hidden layers increases.
(b) Predictions and uncertainty estimates for each method when $L=6$ and Gaussian priors.
Figure 3: Prediction performance of BNNs as the depth increases.
General summary. In terms of the training time, HMC becomes less and less feasible with the increase in depth. With the need to explore high-dimensional parameter spaces, multimodality of the posteriors should be kept in mind as an arising challenge for both mfVI and HMC. In terms of the balance between accuracy and UQ, the mean-field variational inference with ReLU activation function is able to outperform MCMC with the increase in depth.